One eighth of the square root of 7 less than a number is 2. What is the number?
Answer by edjones(8007)

(sqrt(7)/8)-x=2
-x=2-(sqrt(7)/8)
x=(sqrt(7)/8)-2
x=.330719-2
x=-1.66928
.
Ed